$10,000 / 3(y+5)=$333.33 (the slash means division)
What I did so far:
3 times y equals 3y
3 times 5 equals 15
$10,000 / 3y+15=$333.333
I'm not sure what's the next step to take.

Answer by rfer(16322)

from your finish
multiply by (3y+15)
10000=333.33(3y+15)
divide by 333.33
10000/333.33=3y+15
30=3y+15
subtract 15
30-15=3y+15-15
15=3y
divide by 3
5=y
y=5